OpenStudy (anonymous):

Question:Which of the following is the rational exponent expression of 5^√7n? Answers: (7n)^5 7n 1 over 5 5n^7 (7n) 1 over 5

OpenStudy (anonymous):

does this say \[\sqrt[5]{7n}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so then the last one because \[\sqrt[5]{7n}=(7n)^{\frac{1}{5}}\]